# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

### v2

@@ 5 Conclusion
We identified a limitation of standard CFG in counterfactual generation: it applies a uniform guidance weight to all conditioning attributes, leading to amplification of attributes that are meant to remain unchanged. To address this, we proposed Decoupled Classifier-Free Guidance (DCFG), a flexible framework that allows for arbitrary grouping of semantic attributes, with distinct guidance weights applied to each group during sampling, under the mild assumption that the groups are mutually independent. In this work, we instantiate DCFG for counterfactual image generation by separating attributes into two groups: intervened attributes, which include the intervened variables and their descendants in the causal graph, and invariant attributes, which are unaffected by the intervention and expected to remain unchanged. This structure enables DCFG to prevent spurious changes in invariant factors while preserving the intended effect of the intervention. While our focus is on counterfactuals, DCFG can be applied more broadly to conditional generation tasks that require group-wise control over conditioning signals. We evaluated DCFG on CelebA-HQ, EMBED, and MIMIC-CXR, covering natural and medical image domains. Our results show that DCFG significantly reduces attribute amplification while maintaining intervention effectiveness with improved identity preservation, particularly at higher guidance strengths.
Limitations and Future Work. While DCFG demonstrates promising results, the selection of $w_{\text{aff}}$ and $w_{\text{inv}}$ relies on empirical tuning. Future work could explore learned strategies to adaptively tune these weights based on the input condition or diffusion timestep. Additionally, DCFG’s modular nature opens up opportunities for broader integration. One direction is combining DCFG with latent diffusion models or diffusion autoencoders to enable high-resolution synthesis and stronger identity preservation. Another promising extension is applying DCFG selectively at specific diffusion timesteps. Prior work suggests that restricting guidance to mid-to-late timesteps can mitigate over-saturation and improve generation quality. Exploring such selective or dynamic scheduling strategies within the DCFG framework may further enhance counterfactual fidelity.
@@ Acknowledgements
R.R. is supported by the Engineering and Physical Sciences Research Council (EPSRC) through a Doctoral Training Partnerships PhD Scholarship. A.K. is supported by UKRI (grant no. EP/S023356/1), as part of the UKRI Centre for Doctoral Training in Safe and Trusted AI. B.G. received support from the Royal Academy of Engineering as part of his Kheiron/RAEng Research Chair. B.G. and F.R. acknowledge the support of the UKRI AI programme, and the EPSRC, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1).
@@ A.2 Evaluating Counterfactuals
$\tilde{\mathbf{x}}:=\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}})=f_{\theta}(f_{\theta}^{-1}(\mathbf{x},\mathbf{pa}),\tilde{\mathbf{pa}}).$ (20)
We describe three key metrics used to assess counterfactual quality: composition, reversibility, and effectiveness [32, 30].
@@
$\text{Eff}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}):=L_{k}(\tilde{\mathbf{pa}}_{k},\mathbf{Pa}_{k}(\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}))).$ (23)
A note on Composition. We do not report the Composition metric in our evaluation, as it is ill-defined in the context of CFG and DCFG. Since both methods use the same trained diffusion model, applying a null intervention (i.e., $\tilde{\mathbf{pa}}=\mathbf{pa}$) does not meaningfully differentiate between them. If reconstruction is performed without guidance, CFG and DCFG are equivalent—reducing to standard decoding. If guidance is applied, it becomes unclear how to split attributes into invariant ($\mathbf{pa}_{\text{inv}}$) and intervened ($\mathbf{pa}_{\text{aff}}$) groups during null intervention. For instance, assigning all attributes to $\mathbf{pa}_{\text{inv}}$ with $\omega_{\text{inv}}=1$ would effectively disable guidance, making the comparison trivial and uninformative. For this reason, we focus on Effectiveness and Reversibility, which better capture the behavior of guided sampling under interventions.
@@
Reversibility Metrics. We use Mean Absolute Error (MAE) and LPIPS [69] to evaluate reversibility. These metrics quantify the pixel-level and perceptual similarity, respectively, between the original image and its reversed counterfactual.
Appendix B Decoupled Classifier-Free Guidance
We provide a theoretical justification for the Decoupled Classifier-Free Guidance formulation presented in Proposition 1, interpreting it as gradient ascent on a sharpened proxy posterior under a group-level conditional independence assumption.
Proof of Proposition 1.
@@ Appendix B Decoupled Classifier-Free Guidance
For DCFG, the gradient becomes:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})$ $=\nabla\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla\log p(\mathbf{x}_{t})\right),$ (29)
@@
The corresponding implementation interpolates denoising scores per group:
@@
Appendix C Implementation Details
Architecture
@@ Appendix C Implementation Details
We adopt the commonly used U-net backbone [10] for all diffusion models in this work. We modify it to support CFG and DCFG. Each conditioning attribute is projected via a dedicated MLP encoder (section 3.1), and the resulting embeddings are concatenated with the timestep embedding. During training, we apply exponential moving average (EMA) to model weights for improved stability. All images are normalized to the range [-1,1]. The complete architecture and training configurations for each dataset are summarized in table A.2.
@@
Resolution $64\times 64\times 3$ $192\times 192\times 1$ $192\times 192\times 1$
@@
Cond. Embedding Dim $3\times 96$ $2\times 64$ $4\times 64$
@@
Training Procedure for DCFG
The training of our proposed DCFG follows the same setup as standard Classifier-Free Guidance (CFG). Specifically, we apply classifier-free dropout [16] by replacing the entire conditioning vector with a null token (i.e., zero vector) with probability $p_{\varnothing}=0.5$. Unlike the typical choice of $p_{\varnothing}=0.2$, we found that using $p_{\varnothing}=0.5$ better preserves identity, which is particularly important for counterfactual generation tasks. Note that we apply dropout to all attributes jointly, rather than selectively masking subsets. One could alternatively consider group-wise dropout—nullifying only a random subset of attribute groups—but such partial masking may encourage the model to over-rely on the remaining visible attributes, making the resulting guidance less disentangled and less robust. We leave this as an interesting direction for future exploration.
Computation Resources
@@
Appendix D CelebA-HQ
D.1 Dataset details

### v5

@@ 5 Conclusion
We identify a key limitation of classifier-free guidance for counterfactual image generation: using a single global guidance scale for all conditioning attributes can spuriously amplify factors that should remain invariant under a causal intervention. To address this, we propose Factored Classifier-Free Guidance (FCFG), a factored guidance framework that applies separate guidance weights to semantically or causally defined groups of attributes. FCFG supports flexible groupings under a group-wise conditional independence assumption. In this work, we primarily consider a two-group partition that separates affected attributes, i.e., intervened attributes and their descendants, from invariant attributes. We also demonstrate that the framework naturally extends to more fine-grained settings, including multi-attribute and per-attribute configurations. Moreover, FCFG can be integrated with recent CFG refinements such as CFG++ and APG, addressing attribute amplification in these methods as well. We evaluate FCFG on CelebA-HQ, EMBED, and MIMIC-CXR, covering natural and medical image domains. Our results show that FCFG consistently reduces off-target attribute amplification while maintaining intervention effectiveness and improving identity preservation, particularly at higher guidance strengths. The main limitations are that FCFG assumes access to structured conditioning variables and a pre-specified causal or semantic grouping, and that guidance weights are currently chosen empirically. Important directions for future work include adaptively selecting guidance weights based on the input condition or diffusion timestep, and extending factored guidance to settings where the attribute grouping is uncertain or learned. Beyond counterfactual generation, the same principle may also extend to other conditional generation settings and to other diffusion or flow-matching models.
@@ Acknowledgements
T.X. is supported through the Imperial College London UKRI Impact Acceleration Account EP/X52556X/1. B.G. received support from the Royal Academy of Engineering as part of his Kheiron/RAEng Research Chair. B.G. and F.R. acknowledge the support of the UKRI AI programme, and the EPSRC, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1). R.M. is supported by the European Union’s Horizon Europe research and innovation programme under grant agreement 101080302. R.R. is supported by EPSRC through a Doctoral Training Partnerships PhD Scholarship. A.K. is supported by the EPSRC Doctoral Prize.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.”
@@ A.2 Evaluating Counterfactuals
[…]
We describe three key metrics used to assess counterfactual quality: composition, reversibility, and effectiveness (Monteiro et al., 2023; Melistas et al., 2024).
@@
[…]
A note on Composition. We do not report the Composition metric in our evaluation, as it is ill-defined in the context of CFG and FCFG. Since both methods use the same trained diffusion model, applying a null intervention (i.e., $\tilde{\mathbf{pa}}=\mathbf{pa}$) does not meaningfully differentiate between them. If reconstruction is performed without guidance, CFG and FCFG are equivalent, reducing to standard decoding. If guidance is applied, it becomes unclear how to split attributes into invariant ($\mathbf{pa}_{\text{inv}}$) and intervened ($\mathbf{pa}_{\text{aff}}$) groups during null intervention. For instance, assigning all attributes to $\mathbf{pa}_{\text{inv}}$ with $\omega_{\text{inv}}=1$ would effectively disable guidance, making the comparison trivial and uninformative. For this reason, we focus on Effectiveness and Reversibility, which better capture the behavior of guided sampling under interventions.
@@
Reversibility Metrics. We use Mean Absolute Error (MAE) and LPIPS (Zhang et al., 2018) to evaluate reversibility. These metrics quantify the pixel-level and perceptual similarity, respectively, between the original image and its reversed counterfactual.
Appendix B Factored Classifier-Free Guidance
We provide a theoretical justification for the Factored Classifier-Free Guidance formulation presented in Proposition 3.1, interpreting it as gradient ascent on a sharpened proxy posterior under a group-level conditional independence assumption.
Proof of Proposition 3.1.
@@ Appendix B Factored Classifier-Free Guidance
Starting from the group-weighted proxy posterior in Eq. 25, taking logs yields
$\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})=\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\log p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})+\text{const}.$ (26)
Differentiating w.r.t. $\mathbf{x}_{t}$ gives
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t}).$ (27)
To relate $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})$ to conditional scores, we apply Bayes’ rule: $p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})=\frac{p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})p(\mathbf{pa}^{(m)})}{p(\mathbf{x}_{t})}.$ Taking logs and differentiating yields: (noting that $p(\mathbf{pa}^{(m)})$ is constant w.r.t. $\mathbf{x}_{t}$)
$\nabla_{\mathbf{x}_{t}}\log p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}).$ (28)
Substituting Eq. 28 into Eq. 27 gives
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\Big(\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})\Big),$ (29)
@@
From scores to $\boldsymbol{\epsilon}$-parameterization.
In $\boldsymbol{\epsilon}$-prediction diffusion models, scores are implemented by the denoiser $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\cdot)$; we therefore approximate each conditional score by evaluating the model with a masked conditioning input that retains only group $m$. This yields the practical guided noise prediction:
@@
Appendix C DDIM-based Diffusion Counterfactual Generation.
Overview.
Given an observed image $\mathbf{x}_{0}$ and its associated semantic attributes $\mathbf{pa}$, counterfactual generation proceeds as follows: (i) in the abduction step, the observed image is inverted into a latent noise variable via a deterministic DDIM inversion conditioned on the factual attributes; (ii) in the action step, a causal intervention is applied in the semantic attribute space to obtain counterfactual attributes; (iii) in the prediction step, a guided reverse diffusion trajectory is run to generate the counterfactual image. The abduction and action steps are shared across all methods considered in this work, while different guidance formulations affect only the prediction step.
Abduction.
We first describe the abduction step, which aims to infer the latent exogenous noise variable $\mathbf{u}$ corresponding to an observed image $\mathbf{x}_{0}$ under its factual semantic attributes $\mathbf{pa}$. Following prior diffusion-based counterfactual methods (Sanchez and Tsaftaris, 2022; Rasal et al., 2025), we implement abduction via a conditional deterministic DDIM inversion, as summarised in Algorithm 1.
Algorithm 1 Abduction via Deterministic DDIM Inversion (shared across all methods)
Input: observed image $\mathbf{x}_{0}$, factual attributes $\mathbf{pa}$, attribute-split embedder $\{\mathcal{E}_{i}\}_{i=1}^{K}$, DDIM schedule $\{\alpha_{t}\}_{t=0}^{T}$
Embed factual attributes.
[…]
Abduction (DDIM inversion):
1:for $t=0$ to T1T-1 do
2:  $\hat{\mathbf{x}}_{0}\leftarrow\dfrac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right)$
3:  $\mathbf{x}_{t+1}\leftarrow\sqrt{\alpha_{t+1}}\,\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t+1}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$
4:end for
Output: exogenous noise estimate $\mathbf{u}\leftarrow\mathbf{x}_{T}$
Action.
We apply an intervention to the semantic attributes $\mathbf{pa}$ (e.g., do(Male=1)) and propagate its effect through the causal graph using invertible flow-based models, following prior work on flow-based structural causal models (Pawlowski et al., 2020; Ribeiro et al., 2023; Rasal et al., 2025). This produces a counterfactual attribute vector $\widetilde{\mathbf{pa}}$, whose embedding $\tilde{\mathbf{c}}$ is subsequently used for counterfactual prediction. The action step is summarized in Algorithm 2.
Algorithm 2 Action via Attribute Intervention (shared across all methods)
Input: factual attributes $\mathbf{pa}$, intervention index set $\mathcal{I}\subseteq\{1,\dots,K\}$, intervention values $\mathbf{pa}^{\prime}_{\mathcal{I}}$, attribute-split embedder $\{\mathcal{E}_{i}\}_{i=1}^{K}$
Action (Invertible Normalizing Flow-based Structural Causal Models):
[…]
Output: intervened conditioning $\tilde{\mathbf{c}}$
Counterfactual prediction with CFG.
[…]
Algorithm 3 Counterfactual Prediction with CFG
Input: exogenous noise estimate $\mathbf{u}$ (from abduction), counterfactual conditioning $\tilde{\mathbf{c}}$ (from prediction), guidance weight $\omega\geq 0$, DDIM schedule $\{\alpha_{t}\}_{t=0}^{T}$, null conditioning token $\varnothing$
Counterfactual Prediction:
[…]
2:for $t=T$ down to 11 do
3:  ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\epsilon}_{\text{CFG}}}\leftarrow\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)+\omega\big(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)\big)$
4:  $\hat{\mathbf{x}}_{0}\leftarrow\dfrac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\epsilon}_{\text{CFG}}}\right)$
5:  $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\,\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t-1}}\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\epsilon}_{\text{CFG}}}$
6:end for
Output: counterfactual image $\tilde{\mathbf{x}}\leftarrow\mathbf{x}_{0}$
Counterfactual prediction with FCFG.
Counterfactual prediction with FCFG follows the same reverse DDIM decoding procedure as standard CFG, but replaces the global guidance with group-wise guidance defined over masked conditioning embeddings. Given the exogenous noise estimate $\mathbf{u}$ from abduction and the counterfactual conditioning embedding $\tilde{\mathbf{c}}$ from the action step, counterfactual samples are generated by running the deterministic reverse DDIM trajectory using the FCFG-guided score defined in Eq. 14. The full procedure is summarized in Algorithm 4.
Algorithm 4 Counterfactual Prediction with FCFG (Ours; General)
[…]
Counterfactual Prediction:
[…]
2:for $t=T$ down to 11 do
3:  ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\epsilon}_{\mathrm{FCFG}}}\leftarrow\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)$
4:  for $m=1$ to $M$ do
[…]
6:  end for
7:  $\hat{\mathbf{x}}_{0}\leftarrow\dfrac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\epsilon}_{\mathrm{FCFG}}}\right)$
8:  $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\,\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t-1}}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\epsilon}_{\mathrm{FCFG}}}$
9:end for
Output: counterfactual image $\tilde{\mathbf{x}}\leftarrow\mathbf{x}_{0}$
Two-group instantiation.
In most of our experiments, we adopt a two-group partition with affected and invariant attributes (M=2M{=}2), refer to Sec. 3.3 for more details. This can be summarized in Algorithm 5.
Algorithm 5 Counterfactual Prediction with FCFG (Ours; Two-group partition)
[…]
Counterfactual Prediction:
[…]
2:for $t=T$ down to 11 do
[…]
4:  $\hat{\mathbf{x}}_{0}\leftarrow\dfrac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\epsilon}_{\mathrm{FCFG}}}\right)$
5:  $\mathbf{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\,\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t-1}}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\epsilon}_{\mathrm{FCFG}}}$
6:end for
Output: counterfactual image $\tilde{\mathbf{x}}\leftarrow\mathbf{x}_{0}$
Extensions of FCFG.
The core idea of FCFG, namely factoring the guidance signal across causally or semantically defined attribute groups, can be integrated with guidance variants beyond standard classifier-free guidance. In this work, we extend FCFG to CFG++ (Chung et al., 2025) and to adaptive projection guidance (APG) (Sadat et al., 2025) by applying the same factoring mechanism to their respective guided score formulations. The resulting methods, FCFG++ and FAPG, are described in App. J and App. K, respectively, where we present quantitative and qualitative results showing that both CFG++ and APG exhibit attribute amplification, and that integrating FCFG with these methods effectively mitigates this issue.
Appendix D Implementation Details
Architecture.
@@ Appendix D Implementation Details
We adopt the commonly used U-net backbone (Dhariwal and Nichol, 2021) for all diffusion models in this work. We modify it to support CFG and FCFG. Each conditioning attribute is projected via a dedicated MLP embedder (Sec. 3.1), and the resulting embeddings are concatenated with the timestep embedding. During training, we apply exponential moving average (EMA) to model weights for improved stability. All images are normalized to the range [-1,1]. The complete architecture and training configurations for each dataset are summarized in Table A.2.
@@
Resolution $64\times 64\times 3$ $\times 192\times 1 192×192$$\times 192\times 1$
@@
Cond. Embedding Dim $3\times 32$ $2\times 32$ $4\times 32$
@@
Training Procedure for FCFG.
The training of our proposed FCFG follows the same setup as standard Classifier-Free Guidance (CFG). Specifically, we apply classifier-free dropout (Ho and Salimans, 2022) by replacing the entire conditioning vector with a null token (i.e., zero vector) with probability $p_{\varnothing}=0.5$. Unlike the typical choice of $p_{\varnothing}=0.2$, we found that using $p_{\varnothing}=0.5$ better preserves identity, which is particularly important for counterfactual generation tasks. Note that we apply dropout to all attributes jointly, rather than selectively masking subsets. One could alternatively consider group-wise dropout—nullifying only a random subset of attribute groups—but such partial masking may encourage the model to over-rely on the remaining visible attributes, making the resulting guidance less disentangled and less robust. We leave this as an interesting direction for future exploration.
Computation Resources.
@@
Scalability of Attribute-Split Embeddings.
Using a separate MLP for each conditioning attribute may appear unscalable at first glance. However, in our implementation, each per-attribute MLP is intentionally lightweight. Specifically, we use a $1\rightarrow 32\rightarrow 32$ architecture. The resulting 32-dimensional embeddings are concatenated to form a $32\times K$ conditioning vector, where $K$ denotes the number of attributes, and the per-attribute embedding dimension can be adjusted if needed. Even with a relatively large number of attributes (e.g., K=40K{=}40), this design results in only 44,800 parameters in total, which is negligible relative to the overall diffusion model size. In practice, our experiments involve far fewer attributes (typically K=2K{=}233), making the overhead even smaller. By contrast, encoding the same attributes using a single joint MLP that produces a $32\times K$-dimensional conditioning vector would require a substantially larger network. For example, a $40\rightarrow 1280\rightarrow 1280$ MLP contains approximately 1.69 million parameters, making it significantly less parameter-efficient than the attribute-split design. The proposed approach therefore achieves improved parameter efficiency while preserving explicit attribute-level structure in the conditioning space.
Appendix E Additional Baseline and Ablation Results
This appendix provides additional results complementing the main experiments, including comparisons with existing counterfactual generation methods, image-quality evaluation using FID, an ablation on the classifier-free dropout probability $p_{\emptyset}$, and a practical cost comparison. These results further support the conclusion that FCFG improves the target/off-target trade-off while preserving image quality and maintaining a favourable computational profile.
E.1 Comparison with Existing Counterfactual Generation Methods
We compare FCFG with HVAE (Ribeiro et al., 2023), HVAE-soft (Xia et al., 2024), and SA-DCG (Rasal et al., 2025). These methods use structured causal information but incorporate it differently. HVAE and HVAE-soft use predictor-based counterfactual training objectives, while SA-DCG uses a diffusion-autoencoder framework with semantic abduction. In contrast, FCFG modifies only the inference-time guidance rule of a trained conditional diffusion model.
Table A.3 summarizes the comparison. On CelebA-HQ, SA-DCG improves reversibility relative to standard CFG but still exhibits off-target amplification under global guidance. FCFG reduces this amplification while also improving reversibility. On MIMIC-CXR, FCFG achieves comparable or stronger target intervention than HVAE, HVAE-soft, and CFG, while reducing off-target change by a large margin.
Table A.3: Comparison with existing counterfactual generation methods. On CelebA-HQ, results are reported under do(Smiling), with off-target AUC measured on Male. On MIMIC-CXR, off-target AUC denotes the average change over non-intervened attributes. FCFG preserves target intervention effectiveness while substantially reducing off-target amplification. Higher target AUC is better; lower off-target AUC, MAE, and LPIPS are better.
Dataset Method Guidance configuration Target AUC $\uparrow$ Off-target AUC $\downarrow$ MAE $\downarrow$ LPIPS $\downarrow$
CelebA-HQ, do(Smiling) CFG (Ho and Salimans, 2022) $\omega=2.0$ +11.2+11.2 +2.7+2.7 $0.203$ $0.127$
SA-DCG (Rasal et al., 2025) $\omega=2.0$ +12.7+12.7 +3.0+3.0 $0.178$ $0.111$
FCFG (Ours) $\omega_{\mathrm{aff}}=2.0,\omega_{\mathrm{inv}}=1.2$ +10.5+10.5 +0.4+0.4 $\mathbf{0.146}$ $\mathbf{0.098}$
CFG (Ho and Salimans, 2022) $\omega=3.0$ +12.3+12.3 +3.0+3.0 $0.263$ $0.155$
SA-DCG (Rasal et al., 2025) $\omega=3.0$ +12.9+12.9 +3.0+3.0 $0.221$ $0.145$
FCFG (Ours) $\omega_{\mathrm{aff}}=3.0,\omega_{\mathrm{inv}}=1.2$ $\mathbf{+13.1}$ $\mathbf{-1.5}$ $\mathbf{0.177}$ $\mathbf{0.122}$
MIMIC-CXR, do(sex) HVAE (Ribeiro et al., 2023) +7.4+7.4 +9.0+9.0
HVAE-soft (Xia et al., 2024) +7.3+7.3 +5.7+5.7
CFG (Ho and Salimans, 2022) $\omega=3.0$ +7.5+7.5 +10.7+10.7
FCFG (Ours) $\omega_{\mathrm{aff}}=3.0,\omega_{\mathrm{inv}}=1.2$ $\mathbf{+7.6}$ $\mathbf{+0.6}$
MIMIC-CXR, do(finding) HVAE (Ribeiro et al., 2023) +17.1+17.1 +5.1+5.1
HVAE-soft (Xia et al., 2024) +17.3+17.3 +5.0+5.0
CFG (Ho and Salimans, 2022) $\omega=3.0$ +17.8+17.8 +5.3+5.3
FCFG (Ours) $\omega_{\mathrm{aff}}=3.0,\omega_{\mathrm{inv}}=1.2$ $\mathbf{+18.8}$ $\mathbf{+0.6}$
E.2 Image Quality Evaluation
We evaluate general image quality using FID on CelebA-HQ. Although FID is not a complete counterfactual metric, since a valid counterfactual should intentionally differ from the source image along the intervened attribute, it provides a useful complementary measure of realism. As shown in Table A.4, FCFG consistently achieves lower FID than global CFG across all three interventions and guidance strengths. This suggests that reducing off-target amplification also helps preserve overall image quality.
Table A.4: FID comparison on CelebA-HQ. Lower is better. FCFG consistently improves FID over global CFG across interventions.
Method do(Smiling) do(Male) do(Young)
CFG, $\omega=2.0$ $19.0$ $19.3$ $19.8$
FCFG, $\omega_{\mathrm{aff}}=2.0,\omega_{\mathrm{inv}}=1.2$ $\mathbf{18.3}$ $\mathbf{18.8}$ $\mathbf{19.2}$
CFG, $\omega=3.0$ $21.3$ $21.2$ $22.2$
FCFG, $\omega_{\mathrm{aff}}=3.0,\omega_{\mathrm{inv}}=1.2$ $\mathbf{19.6}$ $\mathbf{19.4}$ $\mathbf{20.3}$
E.3 Ablation on Classifier-Free Dropout
Our main experiments use classifier-free dropout probability $p_{\emptyset}=0.5$. To justify this choice, we compare it against the commonly used $p_{\emptyset}=0.2$ on CelebA-HQ under do(Smiling). Table A.5 reports reversibility results for CFG, FCFG, CFG++, FCFG++, APG, and FAPG. Across all guidance formulations, $p_{\emptyset}=0.5$ yields lower MAE and LPIPS than $p_{\emptyset}=0.2$, indicating improved identity preservation. We therefore use $p_{\emptyset}=0.5$ in our main experiments.
Table A.5: Ablation on classifier-free dropout $p_{\emptyset}$ under do(Smiling) on CelebA-HQ. Lower MAE and LPIPS indicate better reversibility and identity preservation. Across all guidance variants, $p_{\emptyset}=0.5$ improves reversibility compared with $p_{\emptyset}=0.2$.
Guidance MAE, $p_{\emptyset}=0.2$ MAE, $p_{\emptyset}=0.5$ LPIPS, $p_{\emptyset}=0.2$ LPIPS, $p_{\emptyset}=0.5$
CFG $0.309$ $\mathbf{0.263}$ $0.180$ $\mathbf{0.155}$
FCFG $0.206$ $\mathbf{0.177}$ $0.140$ $\mathbf{0.122}$
CFG++ $0.298$ $\mathbf{0.251}$ $0.177$ $\mathbf{0.154}$
FCFG++ $0.212$ $\mathbf{0.175}$ $0.149$ $\mathbf{0.121}$
APG $0.280$ $\mathbf{0.269}$ $0.187$ $\mathbf{0.166}$
FAPG $0.231$ $\mathbf{0.195}$ $0.161$ $\mathbf{0.136}$
E.4 Practical Cost and Sampling Overhead
FCFG also offers a favourable simplicity–control trade-off. Unlike HVAE and HVAE-soft, it does not require an additional predictor-based counterfactual training objective or auxiliary pretrained predictors. Unlike SA-DCG, it avoids a heavier diffusion-autoencoder framework and additional identity-preserving inference-time optimization.
The main computational overhead of FCFG occurs at sampling time rather than training time. Standard CFG requires two denoiser evaluations per timestep, one unconditional and one conditional. In contrast, FCFG with $M$ groups requires M+1M{+}1 denoiser evaluations per timestep: one unconditional evaluation and one masked conditional evaluation for each group. In our main two-group setting, this corresponds to three denoiser evaluations per timestep. Thus, the sampling cost increases linearly with the number of guidance groups, while the training-time parameter and memory overhead of the attribute-specific MLP embedders remains negligible relative to the diffusion backbone.
Table A.6 compares FCFG with SA-DCG on CelebA-HQ. FCFG is lighter and faster in this setting while achieving a better target/off-target trade-off and improved reversibility.
Table A.6: Practical comparison with SA-DCG on CelebA-HQ. Results are reported under do(Smiling) with the stronger guidance setting. FCFG is lighter and faster while achieving a better target/off-target trade-off and improved reversibility.
Method Params Time Target AUC $\uparrow$ Off-target AUC $\downarrow$ MAE $\downarrow$ LPIPS $\downarrow$
SA-DCG (Rasal et al., 2025) 72.672.6M $180$s +12.9+12.9 +3.0+3.0 $0.221$ $0.145$
FCFG (Ours) 60.560.5M $140$s $\mathbf{+13.1}$ $\mathbf{-1.5}$ $\mathbf{0.177}$ $\mathbf{0.122}$
Appendix F CelebA-HQ
F.1 Dataset details